A 7.38 g sample of an alloy at 93.0 C is placed into 50.0 g of water at 22.0 C in an insulated coffee cup with a heat capacity of 9.2 J/C. If the final temperature of the system is 31.1 C, what is the specific heat capacity (S Alloy) of the alloy?

A: 4.35 J/g C

B: 1.33 J/g C

C: 0.950 J/g C

D: 3.97 J/g C
Answer #1

eat given by alloy = mass x specific heat x difference in temperature

= 7.38g x sJ/g.K x (31.1-93)K = -456.822 xs J

heat absorbed by water = 50g x 4.18J/g.K x (31.1-22) K =1901.9 J

heat absorbed by calorimeter = heat capacity x difference in temperature

= 9.2J/C x ( 31.1-22) C = 83.72 J

At thermal equilibrium

- heat givenby alloy = heat absorbed by water + calorimeter

-(-456.822s J) = 1901.9 + 83.72

Thus s = 4.3465 J/g.C